Let * be the operation defined by a*b=2b-2a. Find 16*(8*4).
Thanks so much im really confused and i have no idea what to do.

Answer by stanbon(75887)

8*4=2(4)-2(8)=-8
Then 18*(-8)= 2(-8)-2(16)
-16-32=-48
Cheers,
Stan H.